OpenStudy (anonymous):

Write the complex number in standard form. 3-6i/6-i. I came up with 21/37-30/32i. Is this correct?

OpenStudy (amistre64):

conjugate the bottom

OpenStudy (anonymous):

Yep I did . Multiplied by 6+1/6+1

OpenStudy (amistre64):

24 -33i ------- simplify as needed 37

OpenStudy (anonymous):

I came up with 18+3-36i-6i/36-i^2 after multiplying?

OpenStudy (amistre64):

teh bottom part is good; try the top again

OpenStudy (amistre64):

(3-6i) (6+i) 3.6 + 3.i + -6i.6 + -6i.i

OpenStudy (anonymous):

That is still what I am getting?

OpenStudy (amistre64):

24 + (-36+3)i +6

OpenStudy (amistre64):

24 - 33i ; typing and adding just dont mix with me ...

OpenStudy (anonymous):

18+3-36i-6i?

OpenStudy (amistre64):

3.6 + 3.i + -6i.6 + -6i.i what is the answer to this?

OpenStudy (amistre64):

3 times 6 = ?? 3 time i = ?? -6 times 6 = ?? -6i times i = ??

OpenStudy (anonymous):

18+3i-36-6i^2

OpenStudy (amistre64):

good, you even included the typo lol ; lets make that -36i for good measure

OpenStudy (amistre64):

now add like terms

OpenStudy (anonymous):

I think I found my mistake. I took the i and thought it was a 1

OpenStudy (amistre64):

yep; they really shuold use a different symbol for it; something grand and unmistakeable ;)

OpenStudy (anonymous):

ok. I am still having a problem. I went back and re-did the multiplication and came up with a numerator of18-30i-6i(-1) and a denominator of 36-(-1). Is this part correct?

OpenStudy (amistre64):

let see... -36i + 3i ----- -33i the rest is good

OpenStudy (anonymous):

So now I am coming up with 18/37-24/37

OpenStudy (amistre64):

opps; that -6i^2 means -6(-1) = 6

OpenStudy (amistre64):

18 -6i^2 = 18+6 = 24

OpenStudy (anonymous):

isnt i^2=-1?

OpenStudy (amistre64):

yes. that is its definition so the top part is: 18 -33i -6i^2 18-6i^2 -33i 18 -6(-1) -33i 18 +6 -33i 24 -33i

OpenStudy (anonymous):

Ok. and the denominator is -37?

OpenStudy (amistre64):

the denom is: 36 -(i^2) 36 -(-1) 36 +1 37

OpenStudy (anonymous):

Final answer 24/27-33/37i

OpenStudy (anonymous):

oops. Yes 37

OpenStudy (amistre64):

yes; barring typos, that is the correct answer :)

OpenStudy (anonymous):

Woo-Hoo! Thanks for holding my hand through that! I just made a bunch of little mistakes that threw me off.